You are here: Foswiki>AdaptiveOptics Web>Logs>Pisces11Oct013_Log (12 Oct 2013, JuanGuerra)Edit Attach

UT 20131011UT FLAO+PISCES DX

Observer: Guerras,JC(LBTO),Arcidiacono,C, Boutsia, K, (INAF)
Telescope Operator:Becchetti G
SW Support:
AO Support: JChristou
Telescope Support: JLittle
Instrument Support: Edwards,M

Summary

  • Night cancel due to the bad weather.
  • High Humidity
  • Ice on telescope roof.
  • Retro Reflector on all the night to check FLAO.
  • LBTI was on the telescope a bit longer to let them to finish the alignment and the first reconstructors

Details

The observations were cancel due to the bad weather conditions and the accumulation of ice on the roof. The day was mostly cloudy and with temperatures below zero Celcius and with this weather conditions didn't help to meal the ice on the roof and to clear the shutters. The humidity didn't drop for the whole night and some clouds were still passing at the telescope level.

It was decided to keep the RR on the whole night to check FLAO.

A new position for the re-rotators has been determined and the new position for the bin1 is 98.75 and the delta applied to the configuration file is 0.8 degrees.

Filter checks for the flao system + pisces with the ND2.0 installed, it has been found a low correlation of the filter position,..

97.85 ZEBRA RIPPA in AUTOGAIN _130808, _130851, _130924, 67.2, 66.7, 66.4 98.35 ZEBRA 125706 130258 130341 130425 , 64.5, 65.6, 60.2, 60.3, 98.55 131959 132025 65.6, 65.7 98.75 124824 125005 125048 125131 69.5,69.1,69.0, 98.85 ZEBRA 131254 131321 131347 131629 131656 131722 71.4, 63.5, 62.1, 69.1, 69.1, 69.2

-- JuanGuerra - 11 Oct 2013
Topic revision: r2 - 12 Oct 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback